Cutter Enterprises purchased equipment for $84,000 on January 1, 2018. The equipment is expected to have a five-year life and a residual value of $8,700.

Using the sum-of-the-years'-digits method, depreciation for 2018 and book value at December 31, 2018, would be: (Do not round depreciation rate per year)

Multiple Choice

  • $28,000 and $56,000 respectively.

  • $28,000 and $47,300 respectively.

  • $25,100 and $50,200 respectively.

  • $25,100 and $58,900 respectively.
Answer #1

SYD = n(n+1)/2 = 5(5+1)/2 = 15

Depreciation for year 2018

= (84,000-8,700)*5/15

= 75,300*5/15

= 25,100

Book value = Cost - Depreciation = 84,000 - 25,100 = 58,900

Option D is the answer